The period T of a simple pendulum is given by  T=2πLg−−√T=2πLg  where L is the length of the pendulum and g is the acceleration due to gravity. Assume that g = 9.80 m/s2 exactly, and that L, in meters, is lognormal with parameters μL = 0.8 and σ2L=0.05.σL2=0.05.

NOTE: This is a multi-part question. Once an answer is submitted, you will be unable to return to this part.

Find P(T > 3).
Answer #1

ANSWER:-\\\\ Given \,\,that\\\\

P(T>3)\\\\ =P(\frac{2 \pi\sqrt{L}}{\sqrt{g}}>3)\\\\ =P(\sqrt{L}>\frac{3}{2 \pi \sqrt{g}})\\\\ =P(\sqrt{L}>1.4947013)\\\\ =P(L>2.23413)\\\\ =P(ln L >0.80385283)\\\\ =P(Z>\frac{0.80385283-0.8}{\sqrt{0.05}})\\\\ =P(Z>0.01723)\\\\ =0.01723